### v1

@@ Abstract
In recent years, consumer Internet of Things (IoT) devices have become widely used in daily life. With the popularity of devices, related security and privacy risks arise at the same time as they collect user-related data and transmit it to various service providers. Although China accounts for a larger share of the consumer IoT industry, current analyses on consumer IoT device traffic primarily focus on regions such as Europe, the United States, and Australia. Research on China, however, is currently rather rare. This study constructs the first large-scale dataset about consumer IoT device traffic in China. Specifically, we propose a fine-grained traffic collection guidance covering the entire lifecycle of consumer IoT devices, gathering traffic from 70 devices spanning 36 brands and 8 device categories. Based on this dataset, we analyze traffic destinations and encryption practices across different device types during the entire lifecycle and compare the findings with the results of other regions. Compared to other regions, our results show that consumer IoT devices in China rely more on domestic services and overally perform better in terms of encryption practices. However, there are still 20/35 devices improperly conduct certificate validation, and 5/70 devices use insecure encryption protocols. To facilitate future research, we open-source our traffic collection guidance and make our dataset publicly available.
@@ 1 Introduction
While constructing the dataset, we introduce methodological improvements compared to existing studies. Previous studies have primarily focused on the device interaction phase [5, 13]. However, we find differences in traffic across different operational phases of the device and limiting the focus on the interaction phase may lead to an incomplete analysis. Therefore, we propose a fine-grained traffic collection guidance that covers the entire lifecycle of the device to address this limitation. The entire lifecycle refers to all phases during normal device usage, including device setup, user interaction, idle, and deletion [10]. Compared with prior works, the traffic we collected shows greater coverage and can support more fine-grained analysis. Furthermore, we open-source our traffic collection guidance and make our dataset publicly available to facilitate further research.
From March 1 to November 29, 2024, we constructed a dataset of consumer IoT traffic in China based on our traffic collection guidance. Furthermore, for devices that can update firmware during the experiment, we also collected traffic after their firmware update to study the differences in device traffic before and after the firmware update. Based on the traffic collected, we analyzed the traffic destinations and encryption practices and then compared our findings with those of the United States and the United Kingdom [5, 13].
Our study reveals that, regarding traffic destinations, Chinese devices show a higher reliance on domestic services. The distribution of traffic destinations for Chinese devices is mainly directed to China (99.8%) and relies on domestic organizations. In contrast, devices from the United Kingdom have a significantly higher proportion of overseas traffic, with a greater reliance on organizations from the United States. Regarding encryption practices, devices from different regions all show good encryption practices, although some issues remain. In particular, Chinese devices have less PII exposure in unencrypted traffic and a lower percentage of unencrypted traffic compared to devices in the United States and the United Kingdom. However, 20/35 devices still use improper certificate verification, and 5/70 use insecure encryption protocols.
@@
  • The first Chinese consumer IoT devices traffic dataset. We constructed a dataset that include 97.5 hours of traffic from 70 devices across 8 categories and 36 brands. We make this dataset publicly available [^1] https://github.com/NKUHack4FGroup/Lifecycle-Based-Traffic-Dataset to facilitate future research.
  • A lifecycle-based fine-grained traffic collection guidance. We propose a fine-grained traffic collection guidance covering the entire lifecycle of the device in a controlled environment. We open-sourced the guidance and are inviting researchers to build a standard, comprehensive, and state-of-the-art consumer IoT dataset together.
  • Security and privacy measurement of Chinese consumer IoT device traffic. We analyze traffic destinations and encryption practices, then compare the results with those from other regions, revealing the characteristics of Chinese consumer IoT devices and regional differences.
@@ 2 Background
Consumer IoT devices typically refer to IoT devices designed for personal or household use, such as cameras, speakers, and lights. These devices can be bound to and controlled by the companion apps, following similar operational phases (i.e., lifecycle). During these phases, users can control the device through various methods. Meanwhile, consumer IoT devices typically communicate frequently with cloud servers during operation to support various device functions. This section introduces the concepts and definitions relevant to the lifecycle, control methods, and cloud services mentioned above.
@@
Cloud servers of consumer IoT devices. When a consumer IoT device operates, the data it collects is shared with various cloud servers. Following the classification approach used in the study by Ren et al. [5], we categorize these servers into three groups: first party, support party, and third party. First party refers to the manufacturer of the consumer IoT device or the developer of its companion app. Support party refers to organizations that provide computational resources or network services for consumer IoT devices, such as cloud computing, cloud storage service providers, and service providers (e.g., CDN). Third party refers to companies contacted by device and not categorized as first or support parties. We adopt the methodology from Ren et al. [5] to determine server affiliations through a three-step verification process: (1) We first identify the organization of the IP addresses contacted by the device. If the IP’s organization matches the manufacturer or companion app of the consumer IoT device, we classify it as a first party; (2) For unresolved cases, examine official Privacy Policies and Third-Party Data Sharing documents (prioritizing device-specific policies) to identify support party; (3) Categorize any undocumented or unaffiliated servers as the third party.
@@ 3.1 Distribution of Traffic Destinations
We used Chiplot [34] to generate a Sankey diagram (Figure 1) to visualize the distribution of traffic destinations. In terms of destination, 99.8% of all device traffic remained within China, with only 0.2% directed overseas. From the perspective of device types, hubs, humidifiers, lights, sensors, and plugs did not show overseas traffic, likely due to their simple communication with cloud servers. A total of 27 devices showed overseas traffic, all cameras, speakers, or doorbells. For each type of device, the cameras had the highest proportion of foreign traffic at 0.29%, followed by speakers at 0.05% and doorbells at 0.07%.
@@
The United States is the most common overseas destination, accounting for 0.11% of total traffic. In particular, dns.google.com is the most frequently accessed foreign domain, communicated by 20 devices. Other domains involved in United States traffic include ntp.com, amazonaws.com, and iotcplatform.com, supporting services such as time synchronization (e.g., 1.cn.pool.ntp.org), remote access based on TUTK (e.g., all-d-master-tutk.iotcplatform.com, used by Xiaobai doorbells and Jooan cameras), and cloud computing (e.g., p2ps2.coolkit.cn, accessed by Sonoff cameras). Traffic to other countries accounted for less than 0.1% of the total and involved 18 domain names. Among them, 10 domains (e.g., ntp.org, time.cloudflare.com) are related to NTP services. Other services included Google Cloud (e.g., bc.googleusercontent.com), IoT platform management (e.g., all-master.iotcplatform.com), IP address lookup (e.g., api.ipify.org), and AWS cloud hosting (e.g., eu-central-1.compute.amazonaws.com). It should be noted that 27 devices contacted NTP servers in 12 countries and 7 devices contacted NTP servers in multiple countries. This increases the security risk of being affected by international cyber attacks. In contrast, domestic NTP services could offer safer and more stable connections.
@@
To analyze the distribution of traffic across the device lifecycle phases, we visualized the proportional traffic destination during the different lifecycle phases in Figure 2. Since over 99% of the traffic terminates within China, the domestic traffic is omitted from the figure. The results show that during the idle phase, cameras, speakers and doorbells exhibit a relatively high proportion of overseas traffic, and these traffic primarily flow to NTP services. During the setup phase, the cameras and speakers show a notable amount of overseas traffic and involve diverse services, such as DNS service, IP address lookup, IoT platform management, and NTP service. This means the device configurations rely on more types of overseas traffic. During both the interaction and deletion phases, the frequency of overseas destinations decreases. This likely suggests that the main functionalities of IoT devices are built on domestic services.
@@ 3.2 Service Contacted by Devices
Life Cycle Server Camera Doorbell Hub Humidifier Light Plug Sensor Speaker
First Party 3.14 2.33 0.56 2.00 1.33 0.83 1.33 3.29
Setup Support Party 1.96 1.00 0.78 0.00 0.33 0.67 0.00 1.14
Third Party 1.18 0.67 0.00 0.00 0.00 0.00 0.00 0.86
First Party 1.82 1.67 0.44 1.00 1.33 1.08 2.33 2.29
Idle Support Party 2.54 1.00 0.67 0.00 0.33 0.50 0.00 1.57
Third Party 1.29 1.33 0.00 0.00 0.00 0.00 0.33 1.57
First Party 1.43 4.00 0.33 1.00 0.67 0.58 1.33 2.14
Interaction Support Party 1.71 1.33 0.22 0.00 0.17 0.25 0.00 1.71
Third Party 0.93 1.67 0.00 0.00 0.00 0.00 0.33 1.29
First Party 0.43 2.33 0.33 1.00 0.67 0.58 1.00 0.57
Deletion Support Party 0.57 0.00 0.33 0.00 0.17 0.17 0.00 0.29
Third Party 0.21 1.00 0.00 0.00 0.00 0.00 0.33 0.29
First Party 4.18 4.33 1.11 2.00 1.83 1.75 3.67 5.29
Total Support Party 4.89 2.33 1.00 0.00 0.50 0.83 0.00 3.86
Third Party 3.93 2.00 0.00 0.00 0.00 0.00 0.33 3.00
First, regardless of the type of device, first party contacts constituted a significant proportion of all device communications. Second, the interaction with the support party is closely related to the device’s computational resource needs. Among the eight device types, the camera, doorbell, and speaker interact more frequently with the support party. As a result, many manufacturers choose to collaborate with cloud service providers to ensure the devices can leverage cloud computing and storage capabilities, thereby enhancing their stability. The support parties contacted by these devices typically provide cloud computing and storage services (e.g., oss-cn-beijing.aliyuncs.com, ap-beijing.myqcloud.com), video streaming services (e.g., video.qq.com, stream.qqmusic.qq.com), and network communication services (e.g., dns.alidns.com).
Compared with first and support parties, third party domains primarily involve data sharing, advertising services, geolocation, and personalized recommendation services. For example, the Huawei speaker contacts statist.tingmall.com for music streaming services, which may be related to third-party audio components used by the device; the IMOU camera contacts api.weathercn.com for weather query services. Additionally, we also observe that many devices access seemingly unrelated third-party domains. For instance, Xiaomi speaker, Konka camera, Lenovo camera, Pisen camera, and Jooao camera all access the domain www.baidu.com. These observations suggest that while third party services offer various personalized features, they are not explicitly mentioned in privacy policies. As a result, unexpected destination contacts may raise user privacy concerns.
@@
In real-time communication for consumer IoT devices, stable and efficient DNS services are essential. While Google Public DNS is commonly used due to its global network and high availability, GreatBit, as a domestic provider, offers lower latency and better stability for local devices. However, the analysis shows that Google DNS is used more frequently than expected. In consumer IoT device real-time communication scenarios, stable and efficient DNS services are crucial. As a domestic service provider, GreatBit offers lower latency and higher stability for local devices, making it more suitable to meet the demands of domestic devices. Nevertheless, the analysis shows that compared with the local DNS service, the frequency of the device using Google DNS is higher. In addition to Google and Greatbit, the Organizations most frequently contacted by devices include Alibaba Cloud, Xiaomi, and Tencent. The services provided by these companies through their domains include cloud storage services (e.g., oss-cn-shenzhen.aliyuncs.com), NTP services (e.g., time1.cloud.tencent.com, ntp.aliyun.com), and video streaming and media content (e.g., findermp.video.qq.com, vv.video.qq.com), inditcating that devices rely on these companies for the computational resources they provide.
@@
Organization Google Alibaba Cloud Xiaomi Greatbit Tencent Baidu Huawei Tuya NST Jooan
Frequency 20 18 17 17 11 10 9 5 4 3
@@ 3.3 Summary
By analyzing traffic destinations, we find that most Chinese devices rely on domestic services. Specifically, 99.8% of the traffic is directed to domestic servers that depend on domestic organizations. This means that Chinese IoT devices feature a high degree of localization and greater autonomy in control. From the perspective of lifecycle phases, we can infer that the main functionalities of IoT devices, e.g., controlling and monitoring, are built on demostic services. Most foreign services used by Chinese consumer IoT devices are basic Internet services like NTP and DNS, which can also be replaced by Chinese service providers. In addition to contacting first parties, many devices from different manufacturers commonly communicate with major domestic organizations such as Xiaomi, Alibaba, and Tencent. This allows them to receive information from many devices and potentially obtain the capability to construct user profiles, e.g., inferring types of device in a household.

### v2

@@ Abstract
In recent years, consumer Internet of Things (IoT) devices have become widely used in daily life. With the popularity of devices, related security and privacy risks arise at the same time as they collect user-related data and transmit it to various service providers. Although China accounts for a larger share of the consumer IoT industry, current analyses on consumer IoT device traffic primarily focus on regions such as Europe, the United States, and Australia. Research on China, however, is currently relatively rare. This study constructs the first large-scale dataset about consumer IoT device traffic in China. Specifically, we propose a fine-grained traffic collection guidance covering the entire lifecycle of consumer IoT devices, gathering traffic from 77 devices spanning 38 brands and 12 device categories. Based on this dataset, we analyze traffic destinations and encryption practices across different device types during the entire lifecycle and compare the findings with the results of other regions. Compared to other regions, our results show that consumer IoT devices in China rely more on domestic services and overall perform better in terms of encryption practices. However, there are still 23/40 devices improperly conducting certificate validation, and 2/70 devices use insecure encryption protocols. To facilitate future research, we open-source our traffic collection guidance and make our dataset publicly available.
@@ 1 Introduction
While constructing the dataset, we introduce methodological improvements compared to existing studies. Previous studies have primarily focused on the device interaction phase [5, 13]. However, we find differences in traffic across different operational phases of the device, and limiting the focus on the interaction phase may lead to an incomplete analysis. Therefore, we propose a fine-grained traffic collection guidance that covers the entire lifecycle of the device to address this limitation. The entire lifecycle refers to all phases during normal device usage, including device setup, user interaction, idle, and deletion [10]. Compared with prior works, the traffic we collected shows greater coverage and can support more fine-grained analysis. Furthermore, we open-source our traffic collection guidance and make our dataset publicly available to facilitate further research.
According to our traffic collection guidance, we collected the first part of device traffic between March 1 and June 1, 2024, covering 50 consumer IoT devices. To further expand the dataset, we conducted two additional rounds of traffic collection: the second from November 1 to November 29, 2024, and the third from June 7 to June 19, 2025, involving 27 more devices. During the traffic collection experiment, which lasted approximately three months in total, we built a dataset comprising 108 hours of traffic from Chinese consumer IoT devices. The total size of the dataset is 3 GB. Compared with the existing datasets, our dataset covers the full lifecycle traffic of the devices and contains the largest number of devices (within a single region) in the dataset. Furthermore, for devices that can update firmware during the experiment, we also collected traffic after their firmware update to study the differences in device traffic before and after the firmware update. Based on the traffic collected, we analyzed the traffic destinations and encryption practices and then compared our findings with those of the United States and the United Kingdom [5, 13].
Our study reveals that, regarding traffic destinations, Chinese devices show a higher reliance on domestic services. The distribution of traffic destinations for Chinese devices is mainly directed to China (99.75%) and relies on domestic organizations. In contrast, devices from the United Kingdom have a significantly higher proportion of overseas traffic, with a greater reliance on organizations from the United States. Regarding encryption practices, devices from different regions all show good encryption practices, although some issues remain. In particular, Chinese devices have less PII exposure in unencrypted traffic and a lower percentage of unencrypted traffic compared to devices in the United States and the United Kingdom. However, 23/40 devices still use improper certificate verification, and 2/70 use insecure encryption protocols.
@@
  • The first Chinese consumer IoT devices traffic dataset. We constructed a dataset that includes 108 hours of traffic from 77 devices across 12 categories and 38 brands. We have publicly released the dataset111[https://github.com/NKUHack4FGroup/Lifecycle-Based-Traffic-Dataset to support future research within the community.
  • A lifecycle-based fine-grained traffic collection guidance. We propose a fine-grained traffic collection guidance covering the entire lifecycle of the device in a controlled environment. We open-sourced the guidance and invite researchers to build a standard, comprehensive, and state-of-the-art consumer IoT dataset together.
  • Security and privacy measurement of Chinese consumer IoT device traffic. We developed scripts to analyze traffic destinations and encryption practices, then compared the results with those from other regions, revealing the characteristics of Chinese consumer IoT devices and regional differences.
@@ 2 Background
Consumer IoT devices typically refer to IoT devices designed for personal or household use, such as cameras, speakers, and lights. These devices can be bound to and controlled by the companion apps, following similar operational phases (i.e., lifecycle). During these phases, users can control the device through various methods. Meanwhile, consumer IoT devices communicate frequently with cloud servers during operation to support various device functions. This section introduces the concepts and definitions relevant to the lifecycle, control methods, and cloud services mentioned above.
@@
Cloud servers of consumer IoT devices. When a consumer IoT device operates, the data it collects is shared with various cloud servers. Following the classification approach used in the study by Ren et al. [5], we categorize these servers into three groups: first party, support party, and third party. First party refers to the manufacturer of the consumer IoT device or the developer of its companion app. Support party refers to organizations that provide computational resources or network services for consumer IoT devices, such as cloud computing, cloud storage service providers, and service providers (e.g., CDN). Third party refers to companies contacted by the device and not categorized as first or support parties. We adopt the methodology from Ren et al. [5] to determine server affiliations through a three-step verification process: (1) We first identify the organization of the IP addresses contacted by the device. If the IP’s organization matches the manufacturer or companion app of the consumer IoT device, we classify it as a first party; (2) For unresolved cases, examine official Privacy Policies and Third-Party Data Sharing documents (prioritizing device-specific policies) to identify support party; (3) Categorize any undocumented or unaffiliated servers as the third party.
@@ 3.1 Distribution of Traffic Destinations
We used Chiplot [34] to generate a Sankey diagram (Figure 1) to visualize the distribution of traffic destinations. In terms of destination, 99.75% of all device traffic remained within China, with only 0.25% directed overseas. From the perspective of device types, hubs, humidifiers, lights, sensors, and plugs did not show overseas traffic, likely due to their simple communication with cloud servers. A total of 32 devices showed overseas traffic, including cameras, speakers, doorbells, TV, mirror, and pet feeder. For each type of device, the cameras had the highest proportion of foreign traffic at 0.6%, followed by speakers at 0.03%, doorbells at 0.01%, and the other devices at 0.02%.
@@
The United States is the most common overseas destination, accounting for 0.24% of total traffic. In particular, dns.google.com is the most frequently accessed foreign domain, communicated by 21 devices. Other domains involved in United States traffic include ntp.com, amazonaws.com, and iotcplatform.com, supporting services such as time synchronization (e.g., 1.cn.pool.ntp.org), remote access based on TUTK (e.g., all-d-master-tutk.iotcplatform.com, used by Xiaobai doorbells and Jooan cameras), and cloud computing (e.g., p2ps2.coolkit.cn, accessed by Sonoff cameras). Traffic to other countries accounted for less than 0.01% of the total and involved 18 domain names. Among them, 10 domains (e.g., ntp.org, time.cloudflare.com) are related to NTP services. Other services included Google Cloud (e.g., bc.googleusercontent.com), IoT platform management (e.g., all-master.iotcplatform.com), IP address lookup (e.g., api.ipify.org), and AWS cloud hosting (e.g., eu-central-1.compute.amazonaws.com).
From the perspective of cross-border data transmission security, Chinese consumer IoT devices still rely on overseas infrastructure services, which may violate China’s data transmission security regulations. Platforms behind services such as “iotcplatform.com” and “amazonaws.com” depend on complex third party components. This kind of dependency increases the possibility of unauthorized cross-border data flows. Users’ sensitive data may be sent to overseas servers without clear consent, raising the risk of data leaks and misuse. In addition, such cross-border dependencies may affect device availability. For example, network disruptions or policy changes could limit access to remote services, affecting the stability of device functions. Therefore, device manufacturers should consider adopting service localization strategies and reliable domestic infrastructure providers.
It should be noted that 29 devices contacted NTP servers in 12 countries, and 8 devices contacted NTP servers in multiple countries. In contrast, domestic NTP services could offer more stable connections. Therefore, we suggest device manufacturers prioritize the use of domestic NTP sources during product design. This approach reduces potential risks from cross-border access and elevates the overall security protection capabilities of the system.
@@
To analyze the distribution of traffic across the device lifecycle phases, we visualized the proportional traffic destination during the different lifecycle phases in Figure 2. Since over 99% of the traffic terminates within China, the domestic traffic is omitted from the figure. Therefore, the traffic of plug, light, hub, humidifier, and sensor devices is also not included in the figure. The results show that during the idle phase, devices exhibit a relatively high proportion of overseas traffic, and these traffic primarily flow to NTP services. Apart from the doorbell device, during the setup stage, all the devices show a significant amount of overseas traffic and involve diverse services, such as DNS service, IP address lookup, IoT platform management, and NTP service. This means the device configurations rely on more types of overseas traffic. During both the interaction and deletion phases, the frequency of overseas destinations decreases. This likely suggests that the main functionalities of IoT devices are built on domestic services.
@@ 3.2 Service Contacted by Devices
Life Cycle Server Camera Doorbell Hub Humidifier Light Plug Sensor Speaker Other Devices
First Party 3.24 2.33 0.56 2.00 1.33 0.83 1.33 9.22 13.50
Setup Support Party 2.10 1.00 0.78 0.00 0.33 0.67 0.00 0.89 0.00
Third Party 1.34 0.67 0.00 0.00 0.00 0.00 0.00 0.67 6.00
First Party 1.76 1.67 0.44 1.00 1.33 1.08 2.33 9.22 17.75
Idle Support Party 2.48 1.00 0.67 0.00 0.33 0.50 0.00 1.22 1.00
Third Party 1.24 1.33 0.00 0.00 0.00 0.00 0.33 1.22 0.25
First Party 1.38 4.00 0.33 1.00 0.67 0.58 1.33 3.89 11.50
Interaction Support Party 1.69 1.33 0.22 0.00 0.17 0.25 0.00 1.33 0.75
Third Party 0.90 1.67 0.00 0.00 0.00 0.00 0.33 1.00 5.25
First Party 0.41 2.33 0.33 1.00 0.67 0.58 1.00 3.00 8.50
Deletion Support Party 0.59 0.00 0.33 0.00 0.17 0.17 0.00 0.22 0.75
Third Party 0.21 1.00 0.00 0.00 0.00 0.00 0.33 0.22 0.00
First Party 4.03 4.33 1.11 2.00 1.83 1.75 3.67 23.44 35.25
Total Support Party 5.03 2.33 1.00 0.00 0.50 0.83 0.00 3.00 1.25
Third Party 4.00 2.00 0.00 0.00 0.00 0.00 0.33 2.33 7.00
First, regardless of the type of device, first party contacts constituted a significant proportion of all device communications. This indicates that the privacy policies of Chinese devices are relatively well-established. Second, the interaction with the support party is closely related to the device’s computational resource needs. Among the 12 device types, the camera, doorbell, speaker, and mirror interact more frequently with the support party. As a result, many manufacturers choose to collaborate with cloud service providers to ensure the devices can leverage cloud computing and storage capabilities, thereby enhancing their stability. The support parties contacted by these devices typically provide cloud computing and storage services (e.g., oss-cn-beijing.aliyuncs.com, ap-beijing.myqcloud.com), video streaming services (e.g., video.qq.com, stream.qqmusic.qq.com), and network communication services (e.g., dns.alidns.com).
Compared with first and support parties, third party domains primarily involve data sharing, advertising services, geolocation, and personalized recommendation services. For example, the Huawei speaker contacts statist.tingmall.com for music streaming services, which may be related to third party audio components used by the device; the IMOU camera contacts api.weathercn.com for weather query services; the Huawei TV contacts several audio and video service providers, however, we did not find these service providers in their privacy policies. Additionally, we also observe that many devices access seemingly unrelated third party domains. For instance, Xiaomi speaker, Konka camera, Lenovo camera, Pisen camera, and Jooao camera all access the domain www.baidu.com. These observations suggest that while third party services offer various personalized features, they are not explicitly mentioned in privacy policies. As a result, unexpected destination contacts may raise user privacy concerns. Some third party service providers possess strong cross-platform data aggregation capabilities. They can integrate the user behaviors on different devices to build detailed user behavior profiles, which brings serious privacy risks [35, 37]. Therefore, device manufacturers should clearly define data usage boundaries when integrating third party services, transparently disclose relevant third party domains and their purposes in privacy policies, and allow users to make choices.
@@
In real-time communication for consumer IoT devices, stable and efficient DNS services are essential. While Google Public DNS is commonly used due to its global network and high availability, GreatBit, as a domestic provider, offers lower latency and better stability for local devices. However, the analysis shows that Google DNS is used more frequently than expected. In consumer IoT device real-time communication scenarios, stable and efficient DNS services are crucial. As a domestic service provider, GreatBit offers lower latency and higher stability for local devices, making it more suitable to meet the demands of domestic devices. Nevertheless, the analysis shows that compared with the local DNS service, the frequency of the device using Google DNS is higher. In addition to Google and Greatbit, the Organizations most frequently contacted by devices include Alibaba Cloud, Xiaomi, and Tencent. The services provided by these companies through their domains include cloud storage services (e.g., oss-cn-shenzhen.aliyuncs.com), NTP services (e.g., time1.cloud.tencent.com, ntp.aliyun.com), and video streaming and media content (e.g., findermp.video.qq.com, vv.video.qq.com), indicating that devices rely on these companies for the computational resources they provide.
@@
Organization Google Alibaba Cloud Xiaomi Greatbit Baidu Tencent Huawei Tuya NST Jooan
Frequency 21 20 20 17 13 12 10 5 4 3
@@ 3.3 Summary
By analyzing traffic destinations, we find that most Chinese devices rely on domestic services. Specifically, 99.75% of the traffic is directed to domestic servers that depend on domestic organizations. This means that Chinese IoT devices feature a high degree of localization and greater autonomy in control. From the perspective of lifecycle phases, we can infer that the main functionalities of IoT devices, e.g., controlling and monitoring, are built on domestic services. Most foreign services used by Chinese consumer IoT devices are basic Internet services like NTP and DNS, which can also be replaced by Chinese service providers. In addition to contacting first parties, many devices from different manufacturers commonly communicate with major domestic organizations such as Xiaomi, Alibaba, and Tencent. This allows them to receive information from many devices and potentially obtain the capability to construct user profiles, e.g., inferring types of devices in a household.